OpenStudy (anonymous):

A train leaves Boston at 4:00p.m. A second train leaves the same city in the same direction at 6:00p.m. The second train travels 42 mph faster than the first one. IF the second train overtakes the first at 8:00p.m., what is the speed of each of the two trains?

OpenStudy (a_clan):

v1 and v2 be the speeds v2=v1+42 mph Since, they cover same distances in 4 hr and 2 hr respectively t1 * v1 = t2 *v2 4* v1 = 2 * v2 4v1 = 2(v1+42) 2v1= 84 v1 = 42 mph and v2= 84 mph